OpenStudy (anonymous):

A spherical balloon is inflated so that its volume is increasing at the rate of 3.8 ft3/min. How rapidly is the diameter of the balloon increasing when the diameter is 1.8 feet?

OpenStudy (amistre64):

seems like we want a formula for the colume of a sphere to play with

OpenStudy (amistre64):

*volume

OpenStudy (anonymous):

4/3*pi*r^2

geerky42 (geerky42):

^3

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

^3 but, now lets work this for diameter. d = 2r, (2r)^2 = 4r^2

OpenStudy (amistre64):

we dont need to convert, lets just dig in with the derivative ...

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so what is our derivative of the volume equation ... ill prolly work it with the conversion afterwards just for fun tho

OpenStudy (xapproachesinfinity):

Surface area no? that's what we need to differentiate

OpenStudy (anonymous):

dV/dt=4*pi*r^2dr/dt

OpenStudy (amistre64):

we have a rate of change of volume in the information, so lets play with that :) V' = 4 pi r^2 r' yes

OpenStudy (anonymous):

ok so we just sub in 3.8 for r?

OpenStudy (xapproachesinfinity):

the rate of change of the diameter is just the rate of change of the radius

OpenStudy (amistre64):

4r^2 = (2r)^2 = d^2 V' = pi d^2 r' d = 2r d' = 2r' ; r' = d'/2 V' = pi d^2 d'/2

OpenStudy (amistre64):

input the values and solve for d'

OpenStudy (anonymous):

what do i put for v'

OpenStudy (amistre64):

v' is defined as the rate of change of volume ...

OpenStudy (anonymous):

so 3.8

OpenStudy (amistre64):

yep

OpenStudy (xapproachesinfinity):

they gave it to you the volume is increasing with the rate of 3.8

OpenStudy (anonymous):

d'=.747

OpenStudy (amistre64):

V = 4r^3 pi/3 r = d/2 V = 4(d/2)^3 pi/3 V = 4d^3/8) pi/3 V = d^3 pi/6 is our 'adjusted' volume V' = 3d^2 d' pi/6 V' = d^2 d' pi/2 just like before lol

OpenStudy (anonymous):

.747?

OpenStudy (amistre64):

3.8 = (1.8)^2 pi d'/2 2(3.8)/((1.8)^2 pi) = d' = .7466 http://www.wolframalpha.com/input/?i=2%283.8%29%2F%28%281.8%29%5E2+pi%29

OpenStudy (anonymous):

is that the final answer?

OpenStudy (amistre64):

what was it they were looking for?

OpenStudy (anonymous):

the diameter is increasing at ___ ft/min

OpenStudy (amistre64):

then yeah, we determine the rate of change of the diameter

OpenStudy (anonymous):

sweet thanks!!

OpenStudy (amistre64):

yw